OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution of this as an ordered pair. The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution of this as an ordered pair. @Mathematics

OpenStudy (jamesj):

Suppose x is one of the numbers. Then let x + 26 be the other. Then you know that x + (x + 26) = 170. Now solve for x.

OpenStudy (anonymous):

To solve, let's call the numbers random variables first. Let's say number 1 is x and number 2 is y. Translate the given statements to equations: The sum of the two numbers is 170: x+ y = 170 The difference in 26: x-y = 26 If we add the two sides of the equation together we get (x+y) + (x-y) = 170+26 2x = 196 x = 98 We have one number, so take another equation to figure out the other numer: x-y = 26 y = x-26, we know x = 98 y = 98-26 = 72 The two numbers are 98 and 72.

hero (hero):

What the.... I provided a solution for this already...

OpenStudy (anonymous):

I'm confused, I thought they were two different questions.